# Training-Free Bayesianization for Low-Rank Adapters of Large Language Models

### v1

@@ C.1 TFB Variants: $\epsilon$s and Optimal Calibration
Table 8: Performance of different methods applied to LoRA on Llama2-7B pre-trained weights, where Accuracy (ACC) and Expected Calibration Error (ECE) are reported in percentages. The evaluation is done across six common-sense reasoning tasks with a shared hyper-parameter setting after 5,000 gradient steps. We sample $N=10$ during inference in all sampling-based methods including BLoB (Wang et al., 2024d) and TFB variants. Rows with shading indicate training-free Bayesianization methods that use a pre-trained LoRA as their mean. For TFB, the anchor dataset ${\mathcal{D}}$ is set to a randomly sampled subset of the original training set, the performance evaluation metric $l$ is set to accuracy, and the performance drop tolerance is set adaptively to 1% or 0.5% based on whether the given mean overfits. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface and underlining denote the best and the second-best performance, respectively.
@@
LAP 69.20$\pm$1.50 66.78$\pm$0.6911footnotemark: 1 80.05$\pm$0.22 75.55$\pm$0.36 82.12$\pm$0.67 86.95$\pm$0.09 69.14$\pm$1.15 74.94$\pm$0.96 44.10$\pm$1.30 31.60$\pm$0.49
MLE 68.99$\pm$0.58 69.10$\pm$2.84 85.65$\pm$0.92 74.53$\pm$0.66 81.52$\pm$0.25 86.53$\pm$0.28 66.20$\pm$0.87 75.12$\pm$0.85 40.62$\pm$2.25 28.82$\pm$1.30
 + TFB ($\epsilon$=1%) 69.83$\pm$1.02 68.13$\pm$1.03 85.33$\pm$0.97 75.95$\pm$0.34 82.40$\pm$0.20 87.31$\pm$0.26 68.58$\pm$2.77 78.23$\pm$1.20 34.00$\pm$5.57 31.00$\pm$5.00
 + TFB ($\epsilon$=0.5%) 70.89$\pm$0.75 68.58$\pm$1.47 86.21$\pm$0.90 76.27$\pm$0.21 82.80$\pm$0.35 87.66$\pm$0.35 69.93$\pm$2.11 78.87$\pm$1.06 34.67$\pm$3.51 31.00$\pm$2.00
 + TFB* (Optimal ECE) 66.59$\pm$2.13 65.77$\pm$2.44 85.74$\pm$0.63 75.63$\pm$0.48 83.00$\pm$1.31 87.63$\pm$0.23 69.48$\pm$0.85 75.06$\pm$1.34 38.67$\pm$2.52 29.00$\pm$7.81
MAP 68.62$\pm$0.71 67.59$\pm$0.40 86.55$\pm$0.55 75.61$\pm$0.71 81.38$\pm$0.65 86.50$\pm$0.41 69.59$\pm$0.33 75.47$\pm$0.73 44.79$\pm$0.00 28.47$\pm$1.20
 + TFB ($\epsilon$=1%) 69.17$\pm$1.08 67.68$\pm$1.73 84.86$\pm$0.81 75.87$\pm$0.40 81.87$\pm$0.64 86.76$\pm$0.34 68.58$\pm$2.36 78.29$\pm$0.83 33.33$\pm$5.69 32.67$\pm$5.69
 + TFB ($\epsilon$=0.5%) 69.44$\pm$1.42 69.14$\pm$0.70 85.86$\pm$0.37 75.92$\pm$0.39 83.07$\pm$0.61 87.74$\pm$0.23 69.37$\pm$2.54 78.76$\pm$0.87 34.33$\pm$5.51 31.00$\pm$1.00
 + TFB* (Optimal ECE) 69.12$\pm$1.42 66.89$\pm$2.68 86.27$\pm$0.93 75.16$\pm$0.73 82.53$\pm$1.42 87.60$\pm$0.34 69.59$\pm$1.22 75.29$\pm$0.83 37.67$\pm$1.53 29.00$\pm$6.00
 + TFB ($\epsilon$=1%) 69.94$\pm$1.68 70.72$\pm$2.25 85.97$\pm$0.89 73.13$\pm$2.38 83.33$\pm$0.83 85.76$\pm$0.51 70.38$\pm$0.98 79.46$\pm$0.27 40.67$\pm$3.21 32.00$\pm$2.00
 + TFB ($\epsilon$=0.5%) 70.33$\pm$2.01 70.83$\pm$1.56 86.74$\pm$0.97 75.50$\pm$0.23 83.13$\pm$0.76 86.36$\pm$0.26 70.38$\pm$1.03 80.16$\pm$0.71 42.67$\pm$1.15 30.67$\pm$1.53
 + TFB* (Optimal ECE) 68.99$\pm$0.34 69.26$\pm$1.22 86.97$\pm$0.30 75.50$\pm$0.67 82.87$\pm$1.03 87.16$\pm$0.29 70.27$\pm$0.89 77.35$\pm$0.41 37.33$\pm$2.31 30.33$\pm$2.08
LAP 4.15$\pm$1.12 16.25$\pm$2.6111footnotemark: 1 33.29$\pm$0.57 7.40$\pm$0.27 8.70$\pm$1.77 1.30$\pm$0.33 5.84$\pm$0.64 8.51$\pm$1.06 10.76$\pm$3.41 13.91$\pm$0.90
MLE 29.83$\pm$0.58 29.00$\pm$1.97 13.12$\pm$1.39 20.62$\pm$0.74 12.55$\pm$0.46 3.18$\pm$0.09 22.20$\pm$0.39 16.47$\pm$0.86 21.72$\pm$0.30 29.60$\pm$1.29
 + TFB ($\epsilon$=1%) 16.26$\pm$0.36 6.93$\pm$1.43 9.44$\pm$1.28 8.78$\pm$0.84 5.77$\pm$1.37 1.92$\pm$0.77 8.07$\pm$2.77 3.24$\pm$0.76 15.95$\pm$2.54 14.28$\pm$6.16
 + TFB ($\epsilon$=0.5%) 23.32$\pm$1.57 10.93$\pm$1.46 5.82$\pm$0.87 12.43$\pm$0.89 4.60$\pm$0.62 2.30$\pm$0.50 8.47$\pm$2.04 4.64$\pm$0.75 15.87$\pm$5.17 16.77$\pm$4.10
 + TFB* (Optimal ECE) 5.62$\pm$1.05 6.63$\pm$0.97 4.96$\pm$0.83 4.77$\pm$0.59 3.64$\pm$1.33 1.39$\pm$0.34 7.39$\pm$0.95 5.07$\pm$0.10 12.55$\pm$0.95 19.30$\pm$6.49
MAP 29.76$\pm$0.87 29.42$\pm$0.68 12.07$\pm$0.55 23.07$\pm$0.14 13.26$\pm$0.82 3.16$\pm$0.23 19.31$\pm$1.46 15.68$\pm$0.51 17.55$\pm$1.95 30.25$\pm$2.18
 + TFB ($\epsilon$=1%) 11.72$\pm$0.56 6.07$\pm$1.89 11.53$\pm$0.99 5.21$\pm$0.86 6.78$\pm$1.85 4.59$\pm$1.32 7.12$\pm$3.42 4.51$\pm$0.33 14.29$\pm$5.17 14.03$\pm$4.48
 + TFB ($\epsilon$=0.5%) 13.21$\pm$0.30 6.69$\pm$0.52 6.99$\pm$0.96 7.17$\pm$1.28 3.82$\pm$0.60 2.65$\pm$0.30 8.39$\pm$0.75 4.86$\pm$1.03 16.11$\pm$3.22 16.35$\pm$2.94
 + TFB* (Optimal ECE) 9.89$\pm$0.41 6.45$\pm$1.61 3.61$\pm$1.27 4.27$\pm$0.48 2.96$\pm$0.05 1.19$\pm$0.38 5.01$\pm$1.15 4.33$\pm$0.31 12.08$\pm$0.67 19.53$\pm$5.25
 + TFB ($\epsilon$=1%) 6.33$\pm$1.04 5.77$\pm$0.32 10.55$\pm$0.41 4.07$\pm$1.65 4.48$\pm$0.92 7.23$\pm$0.60 6.41$\pm$0.93 4.15$\pm$0.26 10.76$\pm$2.56 17.33$\pm$3.40
 + TFB ($\epsilon$=0.5%) 6.64$\pm$1.39 8.65$\pm$0.28 3.03$\pm$0.43 8.85$\pm$2.42 5.94$\pm$0.46 5.37$\pm$0.44 12.28$\pm$1.24 8.07$\pm$1.01 12.36$\pm$1.73 22.02$\pm$0.30
 + TFB* (Optimal ECE) 4.15$\pm$0.67 4.77$\pm$1.58 2.57$\pm$0.56 4.03$\pm$1.14 3.50$\pm$0.53 1.41$\pm$0.68 8.35$\pm$0.51 5.39$\pm$0.21 14.01$\pm$4.76 17.10$\pm$1.23
LAP 0.60$\pm$0.00 1.03$\pm$0.0411footnotemark: 1 0.88$\pm$0.00 0.57$\pm$0.01 0.52$\pm$0.01 0.31$\pm$0.00 0.81$\pm$0.00 0.70$\pm$0.02 1.35$\pm$0.03 1.36$\pm$0.01
MLE 3.17$\pm$0.37 2.85$\pm$0.27 1.17$\pm$0.13 0.95$\pm$0.07 0.73$\pm$0.03 0.32$\pm$0.00 1.16$\pm$0.00 0.92$\pm$0.03 1.56$\pm$0.06 1.66$\pm$0.05
 + TFB ($\epsilon$=1%) 0.86$\pm$0.06 0.98$\pm$0.02 0.51$\pm$0.04 0.59$\pm$0.01 0.55$\pm$0.02 0.31$\pm$0.00 0.86$\pm$0.03 0.70$\pm$0.05 1.47$\pm$0.02 1.43$\pm$0.06
 + TFB ($\epsilon$=0.5%) 1.58$\pm$0.20 1.03$\pm$0.04 0.48$\pm$0.04 0.65$\pm$0.02 0.54$\pm$0.02 0.30$\pm$0.00 0.87$\pm$0.03 0.70$\pm$0.05 1.46$\pm$0.03 1.43$\pm$0.05
 + TFB* (Optimal ECE) 0.67$\pm$0.00 1.05$\pm$0.06 0.52$\pm$0.01 0.56$\pm$0.01 0.56$\pm$0.02 0.30$\pm$0.00 0.85$\pm$0.03 0.73$\pm$0.04 1.41$\pm$0.04 1.49$\pm$0.02
MAP 2.46$\pm$0.34 2.66$\pm$0.11 0.90$\pm$0.05 1.62$\pm$0.29 0.75$\pm$0.01 0.33$\pm$0.00 1.10$\pm$0.07 0.93$\pm$0.04 1.55$\pm$0.06 1.65$\pm$0.03
 + TFB ($\epsilon$=1%) 0.72$\pm$0.03 0.96$\pm$0.03 0.53$\pm$0.04 0.55$\pm$0.01 0.56$\pm$0.01 0.33$\pm$0.01 0.87$\pm$0.02 0.71$\pm$0.05 1.47$\pm$0.02 1.43$\pm$0.06
 + TFB ($\epsilon$=0.5%) 0.75$\pm$0.04 0.99$\pm$0.02 0.50$\pm$0.04 0.57$\pm$0.01 0.53$\pm$0.02 0.30$\pm$0.00 0.87$\pm$0.02 0.71$\pm$0.05 1.46$\pm$0.02 1.42$\pm$0.05
 + TFB* (Optimal ECE) 0.72$\pm$0.02 1.04$\pm$0.07 0.51$\pm$0.01 0.56$\pm$0.01 0.56$\pm$0.02 0.30$\pm$0.00 0.83$\pm$0.03 0.73$\pm$0.04 1.41$\pm$0.04 1.49$\pm$0.03
 + TFB ($\epsilon$=1%) 0.62$\pm$0.03 0.86$\pm$0.01 0.50$\pm$0.04 0.56$\pm$0.03 0.48$\pm$0.02 0.36$\pm$0.00 0.80$\pm$0.02 0.58$\pm$0.03 1.36$\pm$0.02 1.46$\pm$0.06
 + TFB ($\epsilon$=0.5%) 0.62$\pm$0.03 0.85$\pm$0.01 0.42$\pm$0.03 0.56$\pm$0.01 0.50$\pm$0.01 0.34$\pm$0.00 0.84$\pm$0.03 0.61$\pm$0.01 1.35$\pm$0.01 1.46$\pm$0.06
 + TFB* (Optimal ECE) 0.61$\pm$0.01 0.88$\pm$0.02 0.41$\pm$0.01 0.54$\pm$0.01 0.48$\pm$0.01 0.30$\pm$0.00 0.80$\pm$0.02 0.64$\pm$0.02 1.37$\pm$0.03 1.45$\pm$0.06
C.2 TFB Beyond Low-Rank Isotropic Gaussians
In Table 2 (Sec. 5.3), we compare TFB with two alternative Gaussian distribution families that are controlled by a single parameter $\sigma_{q}$:
  • Constant Standard Deviation Matrix (C-STD): given $\sigma_{q}$, the C-STD’s variational distribution of the weight matrix $q(\operatorname{vec}({\bm{W}}))={\mathcal{N}}(\operatorname{vec}({\bm{W}})|{\bm{\mu}}_{q},{\bm{\Sigma}}_{q})$ where ${\bm{\mu}}_{q}={\bm{W}}_{0}+{\bm{B}}{\bm{A}}$ (same as TFB) and ${\bm{\Sigma}}_{q}=\sigma_{q}^{2}\cdot{\bm{I}}_{n}\otimes[{\bm{B}}{\bm{B}}^{\top}]$.
C-STD’s covariance matrix ${\bm{\Sigma}}_{q}$ is derived through Lemma A.1:
${\bm{\Sigma}}_{q}$ $=[{\bm{I}}_{n}\otimes{\bm{B}}]\cdot[\operatorname{diag}(\operatorname{vec}({\bm{\Omega}})^{2})]\cdot[{\bm{I}}_{n}\otimes{\bm{B}}^{\top}]$ (33)
$=[{\bm{I}}_{n}\otimes{\bm{B}}]\cdot[\sigma_{q}^{2}\cdot{\bm{I}}_{rn}]\cdot[{\bm{I}}_{n}\otimes{\bm{B}}^{\top}]$ (34)
$=\sigma_{q}^{2}\cdot[{\bm{I}}_{n}\otimes{\bm{B}}]\cdot[{\bm{I}}_{n}\otimes{\bm{B}}^{\top}]$ (35)
$=\sigma_{q}^{2}\cdot{\bm{I}}_{n}\otimes[{\bm{B}}{\bm{B}}^{\top}].$ (36)
Table 9 shows the optimal test-set ECE for all three distributions to eliminate the effect of different $\epsilon$ settings. Based on the table, we can see that our low-rank isotropic Gaussian distribution consistently outperforms both FR and C-STD baselines across multiple metrics and datasets. For accuracy (ACC), TFB achieves the highest scores on several datasets, notably reaching 86.97 on ARC-E and 82.87 on OBQA. In terms of calibration, our approach demonstrates superior performance with significantly lower ECE scores across most IND datasets (e.g., 4.15 on WG-S, 4.77 on ARC-C, and 2.57 on ARC-E), often reducing ECE by more than 50% compared to baselines. The method also achieves optimal NLL scores on multiple datasets, including WG-S (0.61) and BoolQ (0.30). These comprehensive results clearly demonstrate that our low-rank isotropic Gaussian distribution not only improves model accuracy but also enhances calibration performance compared to alternative approaches like FR and C-STD.
@@ C.2 TFB Beyond Low-Rank Isotropic Gaussians
Table 9: Ablation study of different training-free Bayesianization methods applied to LoRA on Llama2-7B pre-trained weights, where Accuracy (ACC) and Expected Calibration Error (ECE) are reported in percentages. The numbers reported are based on the $\sigma_{q}^{*}$ for optimal ECE. The evaluation is done across six common-sense reasoning tasks with a shared hyper-parameter setting after 5,000 gradient steps. We sample $N=10$ during inference in all sampling-based methods (BLoB, TFB* (FR), TFB* (C-STD), and TFB* (Ours)). Rows with shading indicate training-free Bayesianization methods that use a pre-trained LoRA as their mean. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface and underlining denote the best and the second-best performance, respectively.

### v2

@@ F.10 Additional Results on Llama2-7B
Table 9: Performance of different methods applied to LoRA on Llama3.1-8B pre-trained weights, where Accuracy (ACC) and Expected Calibration Error (ECE) are reported in percentages. “TF?” denotes whether a method is Training-Free. The evaluation is done across six common-sense reasoning tasks with a shared hyper-parameter setting after fine-tuning of 5 epochs . We sample $N=10$ during inference in all sampling-based methods including BLoB [61] and TFB. Rows with shading indicate training-free Bayesianization methods that use a pre-trained LoRA as their mean. For TFB, we randomly sample a subset of the training data without labels as the anchor dataset ${\mathcal{D}}$. For accuracy-based evaluation ($l_{\text{ACC}}$), we set the performance drop tolerance to $\epsilon=1\%$. For NLL loss ($l_{\text{NLL}}$), we use the same settings as in Table 1. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface and underlining denote the best and the second-best performance, respectively.
@@
ACC ($\uparrow$) MCD 78.03$\pm$0.61 81.64$\pm$1.79 91.37$\pm$0.38 83.18$\pm$0.84 87.20$\pm$1.02 89.93$\pm$0.16 81.42$\pm$1.38 87.27$\pm$0.84 47.92$\pm$2.25 46.53$\pm$0.49
ENS 78.82$\pm$0.52 82.55$\pm$0.42 91.84$\pm$0.36 83.99$\pm$0.74 87.37$\pm$0.67 90.50$\pm$0.14 79.62$\pm$0.57 86.56$\pm$0.60 49.65$\pm$3.22 44.44$\pm$1.96
LAP BP 76.05$\pm$0.92 79.95$\pm$0.42 90.73$\pm$0.08 82.83$\pm$0.85 87.90$\pm$0.20 89.36$\pm$0.52 81.08$\pm$1.20 87.21$\pm$1.20 48.26$\pm$3.93 46.18$\pm$1.30
MLE - 77.87$\pm$0.54 81.08$\pm$0.48 91.67$\pm$0.36 82.30$\pm$0.53 87.90$\pm$0.87 89.58$\pm$0.26 81.48$\pm$2.41 86.83$\pm$0.87 45.83$\pm$0.85 42.36$\pm$1.77
 + TFB ($l_{\text{ACC}}$) 76.40$\pm$0.13 82.00$\pm$0.33 91.39$\pm$0.34 82.37$\pm$0.59 88.07$\pm$0.52 89.66$\pm$0.19 10.66$\pm$1.03 6.44$\pm$1.14 23.59$\pm$1.74 26.90$\pm$4.47
 + TFB ($l_{\text{NLL}}$) 77.44$\pm$0.30 82.53$\pm$1.00 91.33$\pm$0.37 82.53$\pm$0.56 88.53$\pm$0.57 89.75$\pm$0.25 79.76$\pm$1.24 85.52$\pm$0.56 44.33$\pm$4.03 37.00$\pm$2.16
MAP - 76.90$\pm$0.97 81.08$\pm$2.48 91.61$\pm$0.44 82.59$\pm$0.28 85.73$\pm$0.19 90.09$\pm$0.28 79.98$\pm$0.87 86.58$\pm$0.79 43.40$\pm$4.98 38.54$\pm$3.40
 + TFB ($l_{\text{ACC}}$) 76.00$\pm$1.34 82.53$\pm$1.80 91.39$\pm$0.09 82.19$\pm$0.89 86.13$\pm$0.34 89.84$\pm$0.15 79.73$\pm$1.86 86.18$\pm$1.04 42.67$\pm$3.30 36.00$\pm$3.56
 + TFB ($l_{\text{NLL}}$) 76.43$\pm$0.72 82.80$\pm$1.42 91.39$\pm$0.37 82.64$\pm$0.58 86.00$\pm$0.16 89.96$\pm$0.18 80.61$\pm$1.24 86.30$\pm$0.89 45.33$\pm$2.87 35.67$\pm$4.11
BLoB 76.45$\pm$0.37 82.32$\pm$1.15 91.14$\pm$0.54 82.01$\pm$0.56 87.57$\pm$0.21 89.65$\pm$0.15 79.75$\pm$0.43 87.13$\pm$0.00 42.71$\pm$3.71 44.79$\pm$6.64
BLoB-Mean 77.72$\pm$0.12 82.60$\pm$0.60 91.64$\pm$0.55 83.92$\pm$0.48 88.00$\pm$0.80 89.86$\pm$0.05 82.06$\pm$1.15 88.54$\pm$0.31 39.93$\pm$5.20 39.93$\pm$4.02
 + TFB ($l_{\text{ACC}}$) 75.28$\pm$0.33 82.80$\pm$0.33 91.64$\pm$0.15 81.84$\pm$0.74 88.00$\pm$0.16 89.60$\pm$0.35 82.93$\pm$1.91 86.97$\pm$0.62 36.00$\pm$5.66 36.00$\pm$5.72
 + TFB ($l_{\text{NLL}}$) 77.81$\pm$0.36 83.33$\pm$0.19 91.76$\pm$0.48 83.81$\pm$0.39 87.80$\pm$0.16 90.11$\pm$0.28 82.93$\pm$1.54 87.64$\pm$0.51 39.67$\pm$7.32 37.33$\pm$6.65
ECE ($\downarrow$) MCD 16.13$\pm$0.54 13.69$\pm$1.11 6.73$\pm$0.71 13.05$\pm$0.99 9.76$\pm$0.71 7.95$\pm$0.17 13.63$\pm$1.18 9.27$\pm$0.60 30.91$\pm$3.57 33.08$\pm$1.40
ENS 14.72$\pm$0.17 13.45$\pm$1.19 6.59$\pm$0.45 11.17$\pm$0.92 8.17$\pm$0.86 7.35$\pm$0.55 11.37$\pm$1.82 7.21$\pm$1.13 18.92$\pm$6.03 26.80$\pm$3.23
LAP BP 4.18$\pm$0.11 9.26$\pm$3.08 5.27$\pm$0.51 3.50$\pm$0.78 8.93$\pm$0.34 1.93$\pm$0.22 7.83$\pm$1.49 7.80$\pm$1.99 14.49$\pm$0.57 13.17$\pm$2.14
MLE - 17.02$\pm$0.46 16.35$\pm$0.68 7.00$\pm$0.53 13.83$\pm$0.65 9.77$\pm$0.81 8.69$\pm$0.21 14.45$\pm$2.19 10.78$\pm$0.50 32.46$\pm$2.60 38.41$\pm$4.44
 + TFB ($l_{\text{ACC}}$) 8.77$\pm$0.64 9.97$\pm$0.29 4.32$\pm$0.42 6.10$\pm$0.46 5.96$\pm$0.76 5.96$\pm$0.23 10.66$\pm$1.03 6.44$\pm$1.14 23.59$\pm$1.74 26.90$\pm$4.47
 + TFB ($l_{\text{NLL}}$) 12.98$\pm$0.37 11.63$\pm$0.68 5.14$\pm$0.14 10.01$\pm$0.70 7.20$\pm$0.47 7.39$\pm$0.26 6.54$\pm$0.53 5.69$\pm$1.64 14.63$\pm$1.46 19.68$\pm$3.27
MAP - 18.71$\pm$0.74 15.77$\pm$1.60 6.62$\pm$0.64 14.26$\pm$0.92 12.19$\pm$0.55 8.40$\pm$0.25 16.46$\pm$0.44 11.36$\pm$0.58 34.79$\pm$3.76 38.50$\pm$2.18
 + TFB ($l_{\text{ACC}}$) 11.84$\pm$0.98 8.61$\pm$2.53 5.19$\pm$0.43 7.51$\pm$1.92 8.72$\pm$0.75 5.55$\pm$0.64 12.19$\pm$1.63 8.08$\pm$1.24 27.76$\pm$3.44 31.91$\pm$3.68
 + TFB ($l_{\text{NLL}}$) 14.95$\pm$0.65 11.27$\pm$2.53 5.76$\pm$0.63 10.97$\pm$1.19 9.70$\pm$0.69 6.86$\pm$0.31 13.25$\pm$0.95 9.22$\pm$0.91 27.21$\pm$2.62 35.91$\pm$4.12
BLoB 9.93$\pm$0.22 5.41$\pm$1.17 2.70$\pm$0.87 4.28$\pm$0.64 2.91$\pm$0.92 2.58$\pm$0.25 5.61$\pm$0.40 2.48$\pm$0.43 16.67$\pm$0.87 12.78$\pm$4.18
BLoB-Mean 15.43$\pm$0.15 12.41$\pm$1.52 4.91$\pm$0.28 9.37$\pm$1.33 6.44$\pm$0.15 6.26$\pm$0.29 11.22$\pm$0.38 6.34$\pm$0.71 26.65$\pm$3.06 25.40$\pm$5.40
 + TFB ($l_{\text{ACC}}$) 3.04$\pm$0.12 6.76$\pm$1.47 4.81$\pm$1.18 7.42$\pm$1.24 5.26$\pm$0.71 3.22$\pm$0.36 6.55$\pm$1.04 5.54$\pm$1.33 17.00$\pm$4.71 16.65$\pm$4.33
 + TFB ($l_{\text{NLL}}$) 8.16$\pm$0.48 6.48$\pm$0.36 2.44$\pm$0.50 3.83$\pm$0.43 2.67$\pm$0.18 3.10$\pm$0.59 6.69$\pm$1.63 3.61$\pm$0.87 18.45$\pm$6.75 20.53$\pm$6.27
NLL ($\downarrow$) MCD 0.83$\pm$0.01 0.99$\pm$0.10 0.45$\pm$0.06 0.64$\pm$0.03 0.62$\pm$0.08 0.49$\pm$0.01 1.03$\pm$0.02 0.61$\pm$0.03 1.91$\pm$0.18 2.02$\pm$0.15
ENS 0.75$\pm$0.02 0.80$\pm$0.11 0.38$\pm$0.03 0.55$\pm$0.02 0.45$\pm$0.05 0.42$\pm$0.05 0.72$\pm$0.07 0.44$\pm$0.03 1.40$\pm$0.18 1.50$\pm$0.13
LAP BP 0.56$\pm$0.00 1.18$\pm$0.02 1.04$\pm$0.01 0.51$\pm$0.00 0.94$\pm$0.00 0.43$\pm$0.00 1.17$\pm$0.01 1.11$\pm$0.00 1.27$\pm$0.01 1.28$\pm$0.00
MLE - 0.88$\pm$0.04 1.20$\pm$0.11 0.46$\pm$0.04 0.68$\pm$0.01 0.61$\pm$0.06 0.52$\pm$0.01 1.07$\pm$0.06 0.72$\pm$0.06 1.91$\pm$0.16 2.25$\pm$0.21
 + TFB ($l_{\text{ACC}}$) 0.58$\pm$0.02 0.73$\pm$0.02 0.29$\pm$0.02 0.46$\pm$0.01 0.41$\pm$0.03 0.36$\pm$0.00 0.79$\pm$0.04 0.49$\pm$0.05 1.52$\pm$0.09 1.82$\pm$0.09
 + TFB ($l_{\text{NLL}}$) 0.68$\pm$0.03 0.85$\pm$0.02 0.33$\pm$0.03 0.53$\pm$0.01 0.46$\pm$0.04 0.42$\pm$0.00 0.66$\pm$0.02 0.44$\pm$0.01 1.39$\pm$0.11 1.49$\pm$0.05
MAP - 0.99$\pm$0.07 1.12$\pm$0.23 0.46$\pm$0.03 0.74$\pm$0.07 0.79$\pm$0.02 0.52$\pm$0.01 1.19$\pm$0.04 0.83$\pm$0.06 1.97$\pm$0.13 2.32$\pm$0.10
 + TFB ($l_{\text{ACC}}$) 0.65$\pm$0.02 0.65$\pm$0.08 0.35$\pm$0.03 0.48$\pm$0.04 0.54$\pm$0.04 0.35$\pm$0.02 0.88$\pm$0.03 0.60$\pm$0.05 1.58$\pm$0.12 1.95$\pm$0.13
 + TFB ($l_{\text{NLL}}$) 0.77$\pm$0.05 0.80$\pm$0.15 0.38$\pm$0.03 0.57$\pm$0.05 0.61$\pm$0.03 0.40$\pm$0.01 0.96$\pm$0.08 0.66$\pm$0.06 1.69$\pm$0.16 2.12$\pm$0.08
BLoB 0.58$\pm$0.00 0.51$\pm$0.03 0.23$\pm$0.01 0.43$\pm$0.01 0.34$\pm$0.01 0.26$\pm$0.01 0.56$\pm$0.02 0.35$\pm$0.02 1.34$\pm$0.04 1.35$\pm$0.10
BLoB-Mean 0.74$\pm$0.02 0.73$\pm$0.04 0.29$\pm$0.03 0.47$\pm$0.03 0.37$\pm$0.02 0.32$\pm$0.02 0.67$\pm$0.07 0.39$\pm$0.03 1.53$\pm$0.13 1.54$\pm$0.15
 + TFB ($l_{\text{ACC}}$) 0.52$\pm$0.01 0.52$\pm$0.03 0.24$\pm$0.01 0.45$\pm$0.01 0.35$\pm$0.01 0.28$\pm$0.00 0.55$\pm$0.02 0.37$\pm$0.01 1.38$\pm$0.10 1.41$\pm$0.09
 + TFB ($l_{\text{NLL}}$) 0.55$\pm$0.01 0.53$\pm$0.04 0.23$\pm$0.02 0.40$\pm$0.01 0.33$\pm$0.02 0.27$\pm$0.01 0.52$\pm$0.05 0.35$\pm$0.02 1.36$\pm$0.13 1.46$\pm$0.11
@@
Table 10: Performance of different methods applied to LoRA on Llama3.1-8B pre-trained weights, where Accuracy (ACC) and Expected Calibration Error (ECE) are reported in percentages. (${\mathcal{D}}_{\text{train}}$) and (${\mathcal{D}}_{\text{test}}$) denote the anchor dataset is randomly sampled from the training dataset and the testing dataset, respectively. “TF?” denotes whether a method is Training-Free. The evaluation is done across six common-sense reasoning tasks with a shared hyper-parameter setting after fine-tuning of 5 epochs . We sample $N=10$ during inference in all sampling-based methods including BLoB [61] and TFB. Rows with shading indicate training-free Bayesianization methods that use a pre-trained LoRA as their mean. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface and underlining denote the best and the second-best performance, respectively.